# Optimal Real-Time Fusion of Time-Series Data Under Rényi Differential Privacy

### v1

@@ I-A Motivation
Information fusion from multiple sensors is a fundamental technique for environmental perception, monitoring, and control. By integrating heterogeneous measurements, a fused estimate can achieve higher accuracy, robustness, and reliability than any individual sensor alone. However, multi-source data collection and fusion often involve sensitive information, which raises significant privacy concerns. For example, traffic density estimation commonly relies on vehicles’ position and velocity data that can be automatically collected by smart devices, e.g., e.g., GPS-enabled smartphones. Although such measurements are critical for real-time traffic management and congestion prediction, they may inadvertently disclose sensitive personal attributes, such as a driver’s home and workplace locations, thereby exposing private behavioral patterns. To mitigate these risks, data collected from multiple sources are typically aggregated and processed through privacy-preserving mechanisms before being utilized for downstream estimation and control tasks.
@@ I-C Contributions
In this paper, we propose an optimal privacy-aware sensor fusion framework for general nonlinear systems, as illustrated in Fig. 1. The system state $X_{k}$ is observed by $m$ spatially distributed sensors, and the collected measurements are aggregated into the observation vector $Y_{k}=[Y_{k}(i)]_{i=1}^{m}$, where $Y_{k}(i)$ denotes the measurement from the $i$th sensor. To prevent potential privacy leakage caused by direct data disclosure, we design a privacy-aware fusion policy that releases a fused output $Z_{k}$ in an online manner, subject to the constraint that the total privacy leakage over a finite horizon $K$ does not exceed a prescribed privacy budget $\mathsf{B_{G}}$. A remote estimator then utilizes $Z_{k}$ to reconstruct the system state.
@@ IV-A System Model
$Y_{k}=[Y_{k}(i)]_{i=1}^{m},$ (2)
where $Y_{k}(i)\in\mathbb{R}^{n_{y}}$ denotes the local observation obtained from the $i$th sensor. The measurement trajectory $\{Y_{k}\}_{k=1}^{K}$ is private and must be kept confidential. As illustrated in Fig. 1, a fusion policy aggregates the distributed sensor data and releases an instantaneous fused output $Z_{k}\in\mathbb{R}^{n_{z}}$ to a remote estimator. An adversary is assumed to infer the private measurements from the released fusion results, which leads to privacy leakage. Since fusion outputs are released sequentially, the cumulative privacy leakage increases over time. The total allowable privacy leakage over the operation horizon $K$ is referred to as the privacy budget $\mathsf{B_{G}}\in\mathbb{R}^{+}$.
Figure 2: Traffic density estimation through randomly reported position and speed measurements
@@ IV-B Motivating Example
Traffic density on road networks is a key performance indicator for transportation monitoring and control. It can be inferred from randomly collected position and speed measurements reported by smart devices [22]. We consider a scenario in which a subset of vehicles is equipped with similar smart devices, e.g., mobile phones operating under the same operating system or map navigation software, which is common in real-world scenarios. As illustrated in Fig. 2, these devices report GPS position and speed measurements to a centralized fusion center for traffic density estimation. The proposed framework is evaluated using the open-source U.S. Highway 101 dataset [23]. As shown in Fig. 3, the traffic density $X_{k}$ of a highway segment can be accurately estimated from the position and speed measurements $Y_{k}$. While such crowd-sourced sensing enables efficient congestion monitoring, the reported data may inadvertently reveal sensitive information about individual drivers, such as movement patterns or travel habits. To mitigate privacy risks, measurements from each vehicle can be fused by a differentially private fusion sensor prior to traffic density estimation, utilized only over a finite time horizon, and discarded once the associated privacy budget is exhausted.
@@ IV-C Privacy-Aware Fusion and Estimation
At each time step $k$, given the historical measurements $Y^{k}$ and past fusion results $Z^{k-1}$, the stochastic fusion policy processes the sensor measurements according to
@@
where $\mathcal{F}_{k}$ denotes the conditional probability distribution of the fusion output $Z_{k}$. The resulting fused output $Z_{k}$ is then utilized by a remote estimator to produce the state estimate $\tilde{X}_{k}$ according to
@@
To prevent the adversary from inferring the private measurements from the released fusion results, the fusion policy is designed such that the total privacy leakage does not exceed the prescribed privacy budget $\mathsf{B_{G}}$.
@@ IV-D Privacy Measure
$L_{k}\!\left(\mathcal{F}_{k}\left(\cdot\middle|\cdot,Z^{k-1}\right),\alpha\right)=\frac{1}{\alpha-1}\log\sup_{Y^{k},\tilde{Y}^{k}}\mathsf{E}_{\mathcal{F}_{k}\!\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)}\!\left[\left.\left(\frac{\mathcal{F}_{k}\!\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)}{\mathcal{F}_{k}\!\left(Z_{k}\middle|\tilde{Y}^{k},Z^{k-1}\right)}\right)^{\alpha}\right|Z^{k-1}\right],$ (5)
where $Y^{k}$ and $\tilde{Y}^{k}$ denote two adjacent measurement trajectories that differ in exactly one sensor’s measurement, i.e., $Y^{k}(j)=\tilde{Y}^{k}(j)$ for all $j\neq i$ and $Y^{k}(i)\neq\tilde{Y}^{k}(i)$ for some $1\leq i\leq m$. According to [24], the instantaneous fusion mechanism $\mathcal{F}_{k}$ satisfies $\big(\alpha,L_{k}(\mathcal{F}_{k}(\cdot|\cdot,Z^{k-1}),\alpha)\big)$-RDP.
@@
IV-E Optimal Privacy-Aware Fusion Design
Our objective is to develop an optimal fusion strategy and the corresponding state estimator design that minimize the estimation error of system state while ensuring that the total privacy leakage does not exceed the prescribed privacy budget $\mathsf{B_{G}}$. To this end, we formulate the optimal privacy-aware fusion problem as the following constrained optimization problem,
@@ IV-E Optimal Privacy-Aware Fusion Design
$\mathrm{s.t.}\quad$ $\sum_{k=1}^{K}L_{k}\!\left(\mathcal{F}_{k}\!\left(\cdot\middle|\cdot,Z^{k-1}\right),\alpha\right)\leq\mathsf{B_{G}},\quad\forall Z_{k}=z_{k}\in\mathbb{R}^{n_{z}},\;k=1,\ldots,K,$ (7)
where $d\!\left(X_{k},\tilde{X}_{k}\right)$ is a measurable function that quantifies the estimation error at time step $k$. The constraint (7) ensures that the cumulative privacy leakage of each instantaneous funsion mechanism $\mathcal{F}_{k}$, given by $\sum_{k=1}^{K}L_{k}\!\left(\mathcal{F}_{k}\!\left(\cdot\middle|\cdot,Z^{k-1}\right),\alpha\right)$, remains within the prescribed privacy budget $\mathsf{B_{G}}$.
@@ V-A Constrained Optimality Equations
$\mathcal{C}_{k}\left(Z^{k-1}\right)=\left\{\mathcal{F}_{k}\left(z_{k}\middle|y^{k},Z^{k-1}\right),\forall z_{k},y_{j}\in\mathbb{R}^{n_{y}},j=1,2,\cdots k\right\},$ (8)
which includes all conditional distributions that condition on the same historical fusion results $Z^{k-1}$ but different measurements $Y^{k}$. Note that the Rényi privacy loss (5) evaluates all instantaneous fusion polices that has the same conditional information $Z^{k-1}$, i.e., the collection (8). When solving the optimization problem (6)-(7), we cannot obtain the optimal individual sensor fusion policy $\mathcal{F}_{k}\left(\cdot\middle|Y^{k}=y^{k},Z^{k-1}\right)$ when other polices $\mathcal{F}_{k}\left(\cdot\middle|Y^{k}\neq y^{k},Z^{k-1}\right)$ in the collection $\mathcal{C}_{k}$ is unknown. In other words, we need to simultaneously solve all fusion polices with same $Z^{k-1}$, i.e., optimize the policy collection $\mathcal{C}_{k}$.
@@
See in Appendix A. ∎
According to Theorem 1, given the released fusion results $Z^{k-1}$, we can forwardly updated the belief state $b_{k}$ and the remaining privacy $s_{k}$, and backwardly compute the optimal fusion collection by solving the constrained optimality equations (9)-(12). Also, to meet the privacy budget constraint (7), the per-step leakage $L_{k}$ through $\mathcal{C}_{k}$ cannot exceed the remaining privacy budget (12).
@@
Given the information $\left(Y^{k},Z^{k-1}\right)$ and the optimal fusion collection $\mathcal{C}_{k}^{\star}$, the optimal fusion policy $\mathcal{F}_{k}^{\star}\left(\cdot\middle|Y^{k},s_{k},b_{k}\right)$ is obtained by selecting the conditional distribution in $\mathcal{C}_{k}^{\star}$ that corresponds to the actual measurements $Y^{k}$.
As shown in Lemma 1, the optimal fusion policy depends not only on the measurements $Y^{k}$, but also on the remaining privacy budget $s_{k}$ and the belief state $b_{k}$, which distinguishes it from classical DP fusion policies based on the standard composition theorem. The structural difference between the optimal fusion policy and classical fusion policies is illustrated in Fig. 4. As depicted in Fig. 4, the optimal fusion policy adaptively allocates the privacy budget by controlling the remaining budget $s_{k}=s_{k-1}-L_{k-1}$, while simultaneously estimating the system state via the belief state $b_{k}\left(X_{k},\cdot\mid Z^{k-1}\right)$ and controlling the adversary’s belief of the private measurements via $b_{k}\left(\cdot,Y^{k}\mid Z^{k-1}\right)$ in a closed-loop manner.
@@ VI-A Policy Parameterization
$h_{k,y}(i)=\mathcal{H}_{\theta,y}\left(h_{k-1,y},Y_{k}(i)\right),$ (15)
@@
The fusion policy is then parameterized as a conditional Gaussian distribution,
$\mathcal{F}_{\phi,\theta}\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)=\mathcal{N}\Big(Z_{k};\,g_{\phi}(h_{k,z},s_{k})\,f_{\theta}(h_{k,y}),\,I\Big),$ (16)
where $f_{\theta}(h_{k,y})=[f_{\theta}(h_{k,y}(i))]_{i=1}^{m},f_{\theta}(h_{k,y}(i))\in[0,1]^{d}$ is the filtering function applied to each sensor, and $g_{\phi}(h_{k,z},s_{k})\in\mathbb{R}^{m}$ is the fusion vector. Specifically, measurements from different sensors are first processed by the filtering function $f_{\theta}$, then aggregated by the fusion vector $g_{\phi}$, and finally perturbed with Gaussian noise. The structure of the parameterized policies is illustrated in Fig. 5 for clarity.
To sample the fusion result from (16), we first draw Gaussian noise $N_{k}\sim\mathcal{N}(0,I_{d})$ from a standard Gaussian distribution, and then generate $Z_{k}$ via
$Z_{k}=g_{\phi}(h_{k,z},s_{k})f_{\theta}(h_{k,y})+N_{k}.$ (17)
For comparison, most classical differentially private mechanisms are independent of historical released information and use fixed privacy parameters, which can be similarly parameterized as
$Z_{k}=\tilde{g}f_{\theta}(h_{k,y})+N_{k},$ (18)
where the fusion vector $\tilde{g}$ is constant and determined prior to interacting with the measurements. In contrast, our fusion policy adaptively adjusts its output based not only on the current measurements but also on the historical released information, as illustrated in Fig. 5. This structure aligns with the optimal privacy-aware fusion policy described in Theorem 1. We left the further numerical comparisons in Sec. VII to highlight the benefits of using the adaptive fusion vector $g_{\phi}(h_{k,z},s_{k})$.
@@ VI-B Privacy Leakage Evaluation
After parameterization, the privacy leakage of the proposed fusion policy (17) can be efficiently evaluated, significantly reducing the complexity of privacy computation.
@@
VI-C $\left(\alpha,\mathsf{B_{G}}\right)$-RDP Sensor Fusion
Motivated by Theorem 2, we implement the sensor fusion algorithm (Algorithm 1) by embedding the hard constraint (12) via the clipping operation (22). Since the privacy budget constraint is never violated, Algorithm 1 satisfies $\left(\alpha,\mathsf{B_{G}}\right)$-RDP, ensuring differential privacy at the trajectory level [19].
Furthermore, Algorithm 1 can be used to evaluate the parameterized fusion policy $\mathcal{F}_{\theta,\phi}$ and the state estimator $\mathcal{E}_{\omega}$, which is necessary for applying gradient descent techniques to update their parameters. In the following, we describe in detail how to optimize the fusion policy and state estimator based on Algorithm 1.
Algorithm 1 $\left(\alpha,\mathsf{B_{G}}\right)$-RDP sensor fusion
Initialize $k=1$, $s_{1}=\mathsf{B_{G}}$, $Z_{0}=0$, $\mathsf{B}_{0}=0$, $h_{k,y}$, $h_{k,z}$ and $\tilde{h}_{k,z}$.
while $k\leq K$ do
  Collect the measurement $Y_{k}$ from the environment
  Update the recurrent representations via $h_{k,y}$, $h_{k,z}$ and $\tilde{h}_{k,z}$
  Compute the fusion vector $g_{\phi}\big(h_{k,z},s_{k}\big)$ and clip it using (22).
  Release the fusion result: $Z_{k}=g_{\phi}\left(h_{k,z},\mathsf{B}_{k}\right)f_{{\theta}}\left(h_{k,y}\right)+N_{k}$, $N_{k}\sim\mathcal{N}\left(0,I_{d}\right)$
  Compute the state estimate $\mathcal{E}_{\omega}\left(\tilde{h}_{k,z}\right)$
  Update the remaining privacy budget $s_{k+1}=s_{k}-\frac{\alpha d}{2}\bar{g}_{\phi}^{2}\left(h_{k,z},\theta\right)$
  The environment returns the estimation error $d\left(X_{k},\tilde{X}_{k}\right)$
  $k=k+1$
end while

### v2

@@ I-A Motivation
Information fusion from multiple sensors is a fundamental technique for environmental perception, monitoring, and control. By integrating heterogeneous measurements, a fused estimate can achieve higher accuracy, robustness, and reliability than any individual sensor alone. However, multi-source data collection and fusion often involve sensitive information, which raises significant privacy concerns. For example, traffic density estimation commonly relies on vehicles’ position and velocity data that can be automatically collected by smart devices, e.g., GPS-enabled smartphones. Although such measurements are critical for real-time traffic management and congestion prediction, they may inadvertently disclose sensitive personal attributes, such as a driver’s home and workplace locations, thereby exposing private behavioral patterns. To mitigate these risks, data collected from multiple sources are typically aggregated and processed through privacy-preserving mechanisms before being utilized for downstream estimation and control tasks.
@@ I-C Contributions
In this paper, we propose an optimal privacy-aware sensor fusion framework for general nonlinear systems, as illustrated in Fig. 1. The system state $X_{k}$ is observed by $m$ spatially distributed sensors, and the resulting measurements are collected in the observation vector $Y_{k}=[Y_{k}(i)]_{i=1}^{m}$, where $Y_{k}(i)$ denotes the measurement from the $i$th sensor. These measurements are processed by a fusion center to generate the fused output $Z_{k}$, which is then used by a remote estimator to reconstruct the system state. However, the fusion center’s output may also be exploited by an adversary to infer private information, thereby causing privacy leakage. The maximum allowable privacy leakage over the operational horizon $K$ is referred to as the privacy budget, denoted by $\mathsf{B_{G}}\in\mathbb{R}^{+}$. To mitigate privacy leakage arising from direct data disclosure, we design a privacy-aware fusion policy that releases $Z_{k}$ online while ensuring that the total privacy leakage does not exceed the budget $\mathsf{B_{G}}$.
@@ IV-A System Model
$Y_{k}=[Y_{k}(i)]_{i=1}^{m}\sim p\left(Y_{k}\middle|X_{k}\right)$ (2)
where $Y_{k}(i)\in\mathbb{R}^{n_{y}}$ denotes the local observation obtained from the $i$th sensor, which are correlated with the underlying process $X_{k}$. The measurement trajectory $\{Y_{k}\}_{k=1}^{K}$ is private and must be kept confidential. As illustrated in Fig. 1, a fusion policy aggregates the distributed sensory data and releases an instantaneous fused output $Z_{k}\in\mathbb{R}^{n_{z}}$ to a remote estimator, which aims at inferring $X_{k}$ for system monitoring and control. The output of the fusion center might also be used by an adversary to infer private information, which may lead to privacy leakage.
Figure 2: Traffic density estimation through position and speed measurements of vehicles
@@ IV-B Motivating Example
Traffic density on road networks is a key performance indicator for monitoring and control of transportation networks, which can be inferred using the position and speed measurements reported by smart devices, e.g., cell phones [22]. As illustrated in Fig. 2, we consider a motivating scenario in which a subset of vehicles is equipped with similar smart devices that report their real-time GPS-based speeds and positions to a centralized fusion center for traffic density estimation. Fig. 3 shows the estimated traffic density $\left(X_{k}\right)$ for a highway segment can be accurately estimated from the position and speed measurements $\left(Y_{k}\right)$. While this crowd-sourced sensing technique enables efficient congestion monitoring, the reported data may inadvertently reveal sensitive information about individual drivers, such as movement patterns or travel habits. To mitigate privacy risks, measurements from vehicles would be fused by a differentially private fusion technique prior to traffic density estimation.
@@ IV-C Privacy-Aware Fusion and Estimation
At each time step $k$, given the historical measurements $Y^{k}=\left[Y_{1},Y_{2},\cdots Y^{k}\right]$ and past fusion results $Z^{k-1}=\left[Z_{1},Z_{2},\cdots Z^{k-1}\right]$, the stochastic fusion policy processes the sensor measurements according to
@@
where $\mathcal{F}_{k}$ denotes the conditional probability distribution of the fusion output $Z_{k}$. The output of the fusion policy $Z_{k}$ is then utilized by a remote estimator to produce the state estimate $\tilde{X}_{k}$ according to
@@
To prevent the adversary from inferring the private measurements from the released fusion results, the fusion policy is designed such that the total privacy leakage does not exceed the prescribed privacy budget $\mathsf{B_{G}}\in\mathbb{R}^{+}$.
@@ IV-D Privacy Measure
$L_{k}\!\left(\mathcal{F}_{k},Z^{k-1};\alpha\right)=\frac{1}{\alpha-1}\log\sup_{Y^{k},\tilde{Y}^{k}}\mathsf{E}_{\mathcal{F}_{k}\!\left(Z_{k}\middle|\tilde{Y}^{k},Z^{k-1}\right)}\!\left[\left.\left(\frac{\mathcal{F}_{k}\!\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)}{\mathcal{F}_{k}\!\left(Z_{k}\middle|\tilde{Y}^{k},Z^{k-1}\right)}\right)^{\alpha}\right|Z^{k-1}\right],$ (5)
where $Y^{k}$ and $\tilde{Y}^{k}$ denote two adjacent measurement trajectories that differ in exactly one sensor’s measurement, i.e., $Y^{k}(j)=\tilde{Y}^{k}(j)$ for all $j\neq i$ and $Y^{k}(i)\neq\tilde{Y}^{k}(i)$ for some $1\leq i\leq m$. According to [23], the instantaneous fusion mechanism $\mathcal{F}_{k}$ satisfies $\big(\alpha,L_{k}\!\left(\mathcal{F}_{k},Z^{k-1};\alpha\right)\big)$-RDP.
@@
IV-E Optimal Privacy-Aware Fusion Policy Design
Our objective is to develop an optimal fusion policy and the corresponding estimation policy that minimize the estimation error of system state while ensuring that the total privacy leakage does not exceed the prescribed privacy budget $\mathsf{B_{G}}$. To this end, we formulate the optimal privacy-aware fusion problem as the following constrained optimization problem,
@@ IV-E Optimal Privacy-Aware Fusion Policy Design
$\mathrm{s.t.}\quad$ $\sum_{k=1}^{K}L_{k}\!\left(\mathcal{F}_{k},Z^{k-1};\alpha\right)\leq\mathsf{B_{G}},\quad\forall Z_{k}=z_{k}\in\mathbb{R}^{n_{z}},\;k=1,\ldots,K,$ (7)
where $d\!\left(X_{k},\tilde{X}_{k}\right)$ is a measurable function that quantifies the estimation error at time step $k$. The constraint (7) ensures that the cumulative privacy leakage of each instantaneous funsion mechanism $\mathcal{F}_{k}$, given by $\sum_{k=1}^{K}L_{k}\!\left(\mathcal{F}_{k},Z^{k-1};\alpha\right)$, remains within the privacy budget $\mathsf{B_{G}}$.
@@ V-A Constrained Optimality Equations
$\mathcal{C}_{k}\left(Z^{k-1}\right)=\left\{\mathcal{F}_{k}\left(z_{k}\middle|y^{k},Z^{k-1}\right),\forall z_{k}\in\mathbb{R}^{n_{z}},y_{j}\in\mathbb{R}^{m\times n_{y}},j=1,2,\cdots k\right\},$ (8)
which includes all conditional distributions that condition on the historical fusion results $Z^{k-1}$ and different measurements $Y^{k}=y^{k}$. Note that the Rényi privacy loss (5) evaluates all instantaneous fusion polices that has the same conditional information $Z^{k-1}$, i.e., the collection (8). Thus, the optimization problem (6)-(7) cannot be solved by obtaining the optimal individual sensor fusion policy $\mathcal{F}_{k}\left(\cdot\middle|Y^{k}=y^{k},Z^{k-1}\right)$ when other polices $\mathcal{F}_{k}\left(\cdot\middle|Y^{k}\neq y^{k},Z^{k-1}\right)$ in the collection $\mathcal{C}_{k}$ are unknown. In other words, we need to simultaneously solve all fusion polices with same $Z^{k-1}$, i.e., optimize the policy collection $\mathcal{C}_{k}$.
@@
See Appendix A. ∎
According to Theorem 1, given the released fusion results $Z^{k-1}$, we can updated the belief state $b_{k}$ and the remaining privacy $s_{k}$, and compute the optimal fusion collection by solving the constrained optimality equations (9)-(12). Also, to meet the privacy budget constraint (7), the per-step leakage $L_{k}\left(\mathcal{C}_{k};\alpha\right)$ cannot exceed the remaining privacy budget (12).
@@
Given the information $\left(Y^{k}=y^{k},Z^{k-1}\right)$ and the optimal fusion collection $\mathcal{C}_{k}^{\star}$, the optimal fusion policy $\mathcal{F}_{k}^{\star}\left(\cdot\middle|y^{k},s_{k},b_{k}\right)$ is obtained by selecting the conditional distribution in $\mathcal{C}_{k}^{\star}$ that corresponds to the actual measurements $Y^{k}=y^{k}$.
Proof.
As discussed previously, the optimal fusion policy, denoted by $\mathcal{F}\left(\cdot\middle|Y^{k}=y^{k},Z^{k}\right)$, belongs to the optimal fusion collection $\mathcal{C}_{k}^{\star}\left(Z^{k-1}\right)$. By Theorem 1, $\mathcal{C}_{k}^{\star}$ depends only on $s_{k}$ and $b_{k}$, i.e., $\mathcal{C}_{k}^{\star}\left(s_{k},b_{k}\right)$, where $s_{k}$ and $b_{k}$ are computed based on $Z^{k-1}$. Therefore, the optimal fusion policy can also be expressed as a function of $s_{k}$ and $b_{k}$, namely $\mathcal{F}\left(\cdot\middle|Y^{k}=y^{k},s_{k},b_{k}\right)$, and is selected from $\mathcal{C}_{k}^{\star}\left(s_{k},b_{k}\right)$ according to the actual measurements $Y^{k}=y^{k}$. ∎
As established in Lemma 1, the optimal fusion policy depends not only on the measurements $Y^{k}$, but also on the remaining privacy budget $s_{k}$ and the belief state $b_{k}$. This feature distinguishes it from classical differentially private fusion policies based on the standard composition theorem. The structural distinction between these two classes of policies is illustrated in Fig. 4. In particular, the optimal fusion policy allocates the privacy budget adaptively in a closed-loop manner based on $b_{k}$ and $Y_{k}$. Moreover, conditioned on $Z^{k-1}$, the adversary can form its belief about the private measurements through the inference distribution $p\left(Y^{k}\middle|Z^{k-1}\right)$, which is the marginal of the belief state $b_{k}\left(X_{k},Y^{k}\right)=p\left(X_{k},Y^{k}\middle|Z^{k-1}\right)$. According to the belief state update rule, the evolution of the adversary’s belief state is governed by the optimal fusion policy.
@@ VI-A Policy Parameterization
$h_{k,y}(i)=\mathcal{H}_{\theta,y}\left(h_{k-1,y}(i),Y_{k}(i)\right),$ (15)
@@
The fusion policy is parameterized as a conditional Gaussian distribution:
$\mathcal{F}_{\phi,\theta}\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)=\mathcal{N}\Big(Z_{k};,g_{\phi}^{\top}(h_{k,z},s_{k})f_{\theta}(h_{k,y}),I\Big),$ (16)
where $f_{\theta}(h_{k,y})=[f_{\theta}(h_{k,y}(i),i)]_{i=1}^{m}\in[0,1]^{m\times d}$ denotes the filtering function applied to each sensor, with each row $f_{\theta}(h_{k,y}(i),i)$ representing a $d$-dimensional feature vector extracted from a different sensor. Moreover, $g_{\phi}(h_{k,z},s_{k})\in\mathbb{R}^{m}$ denotes the fusion vector with the $i$th element $g_{\phi}(h_{k,z},s_{k},i)$ computed based on $h_{k,z}$ and $s_{k}$. For clarity, the structure of the parameterized policy is illustrated in Fig. 5. Specifically, measurements from different sensors are first processed by the filtering function $f_{\theta}$, then aggregated through the fusion vector $g_{\phi}$, and finally perturbed by Gaussian noise. The proposed fusion policy in (16) is inspired by the differentially private filter structures in [10, 15], which have been shown to achieve a better privacy–utility trade-off compared with schemes that inject noise prior to fusion.
To sample the fusion result from (16), we first draw Gaussian noise $N_{k}\sim\mathcal{N}(0,I)$ from a standard Gaussian distribution, and then generate $Z_{k}$ via
$Z_{k}=g_{\phi}^{\top}(h_{k,z},s_{k})f_{\theta}(h_{k,y})+N_{k}.$ (17)
In contrast, most classical differentially private mechanisms are independent of the history of the released information and use fixed privacy parameters, which can be similarly parameterized as
$Z_{k}=\tilde{g}^{\top}f_{\theta}(h_{k,y})+N_{k},$ (18)
where the fusion vector $\tilde{g}$ is constant which is computed a priori. Differently, our fusion policy adaptively adjusts its output based not only on the current measurements but also on the history of the released information, as illustrated in Fig. 5. This structure aligns with the optimal privacy-aware fusion policy described in Theorem 1.
@@ VI-B Privacy Leakage Evaluation
Evaluating the privacy metric in (5) is computationally expensive for general fusion policies, as it involves divergence calculations over all possible adjacent measurement trajectories. In the following, however, we show that the privacy leakage of the fusion policy admits an analytical expression if it is parameterized as in (16), thereby significantly reducing the computational complexity of policy optimization.
@@
VI-C Sensor Fusion and State Estimation Under the Proposed Scheme
Motivated by Theorem 2, we implement the sensor fusion and state estimation algorithm summarized in Algorithm 1. In particular, the fusion center executes Steps 3–6 to produce the fusion output $Z_{k}$ and enforce the hard privacy budget constraint in (12) via the clipping operation in (22). After receiving $Z_{k}$, the estimator executes Step 8 to estimate the state. Since the privacy budget constraint is strictly enforced at all times, Algorithm 1 satisfies $\left(\alpha,\mathsf{B_{G}}\right)$-RDP and thus guarantees trajectory-level differential privacy [19].
Furthermore, Algorithm 1 can be used to evaluate the parameterized fusion policy $\mathcal{F}_{\phi,\theta}$ and the state estimator $\mathcal{E}_{\omega}$, which is necessary for applying gradient descent techniques to update their parameters. In the following, we describe in detail how to optimize the fusion policy and state estimator based on Algorithm 1.
Algorithm 1 $\left(\alpha,\mathsf{B_{G}}\right)$-RDP sensor fusion and state estimation
1:Initialize $k=1$, $s_{1}=\mathsf{B_{G}}$, $Z_{0}=0$, $\mathsf{B}_{0}=0$, $h_{k,y}$, $h_{k,z}$ and $\tilde{h}_{k,z}$.
2:while $k\leq K$ do
3:  Collect the measurement $Y_{k}$
4:  Update the recurrent representations $h_{k,y}$ and $h_{k,z}$
5:  Compute the fusion vector $g_{\phi}\big(h_{k,z},s_{k}\big)$ and clip it using (22)
6:  Release the fusion result: $Z_{k}=g_{\phi}^{\top}\left(h_{k,z},\mathsf{B}_{k}\right)f_{{\theta}}\left(h_{k,y}\right)+N_{k}$, $N_{k}\sim\mathcal{N}\left(0,I\right)$
7:  Update the remaining privacy budget $s_{k+1}=s_{k}-\frac{\alpha d}{2}\bar{g}_{\phi}^{2}\left(h_{k,z},\theta\right)$
8:  Update the recurrent representation $\tilde{h}_{k,z}$ and compute the state estimate $\mathcal{E}_{\omega}\left(\tilde{h}_{k,z}\right)$
9:  The environment returns the estimation error $d\left(X_{k},\tilde{X}_{k}\right)$
10:  $k=k+1$
11:end while